# Assessment of Simulation-based Inference Methods for Stochastic Compartmental Models

### v2

@@ 2.3 Observation Model
The course of an epidemic is usually only partially observed and data is often only available in aggregated form on population level. For both models introduced above, we consider an observation model reflecting two often measured quantities to monitor infectious diseases. For the SIR model, the observables are
@@
where the first entry is the ratio of currently infected people describing the count data from testing, and the second entry describes seroprevalence, the ratio of individuals with antibodies present due to a prior infection. Because the observation process may have several limitations, an additional scaling parameter $s$ is used to capture that only a fraction of cases is reported [ethiopian_data].
@@
where we truncate $\overline{y}_{1}(t_{i})$ and $\overline{y}_{2}(t_{i})$ at a lower bound of 0 to avoid unrealistic negative observations. As in the dataset given by [ethiopian_data], the standard deviations $\sigma_{i,1},\sigma_{i,2}$ were computed directly using a normal approximation to the binomially distributed testing count data. Alternative probability distributions may be more appropriate in other contexts. Further details on the computations of measurement noise are available in Supplementary Information A.2.
From each of the two stochastic compartmental models we generated synthetical data emulating the time course of a disease outbreak. We generated dense datasets with observations every seven or ten days. Additionally, to create realistic, heterogeneous observation schedules, we randomly subsampled the dense time series into sparse datasets reducing the number of time points from 57 to fewer than 15 per observable. We also introduced partial missingness in the sparse datasets, where only one of the two observables was available at certain time points. Details of the data generation and different datasets used are outlined in Supplementary Information A.6.
@@ 3 Bayesian Inference for Stochastic Compartmental Models
To infer the unknown model parameters $\mathbf{\theta}$, which best explain the data $\mathcal{D}^{*}=\{\overline{y}_{t_{i}}\mid i\in\{1,...,n\}\}$}, we apply Bayesian inference techniques. For this, we denote the space of unobserved states of the diffusion process by $\mathcal{X}$.
@@
The posterior represents the updated belief about the parameters after taking the observed data into account via the likelihood, effectively integrating both previous knowledge and the new information provided by the data. While the full posterior distribution is typically of primary interest, we additionally consider the mode (called maximum a-posteriori) of the posterior distribution as the best point estimate [bernardo1994bayesian].
@@
In most cases, evaluating the density of hidden states analytically is impossible [Doucet_SMC], which prohibits evaluating the likelihood. Even in the case of a tractable likelihood, normalization of the posterior is done via the marginal likelihood, which can remain intractable. High‑dimensional parameter spaces further make discretization or numerical integration prohibitively expensive [wilkinson2010, hasenauer2013thesis].
Approximation techniques that estimate the posterior distribution without directly computing the complicated likelihood function enable robust parameter estimation, even in settings where traditional methods may fail due to mathematical or computational complexity. In this section, we will introduce two approaches (particle filters [golightly_particle_2011] and conditional normalizing flows [cranmer2020frontier]) that circumvent the computation of the likelihood function. PF approximate the likelihood with an unbiased estimator obtained by sequential importance sampling, and conditional normalizing flows learn a series of invertible mappings to transform a Gaussian density into the desired density.
@@ 3.1 Particle Markov Chain Monte Carlo Methods
Particle Markov chain Monte Carlo (PMCMC) methods are based on the idea of using an unbiased estimator of the likelihood function, obtained by a particle filter, inside a Markov chain Monte Carlo (MCMC) method [golightly_particle_2011].
Particle filters are sequential Monte Carlo (SMC) techniques based on importance sampling with resampling used for handling non-linear and non-Gaussian time-series models. Let $X_{t_{0}:t_{M}}$ denote a discrete time-series of the unobserved states and $y_{t_{0}:t_{M}}$ a time-series of observations at the same time points. The main idea is to iterate over the following three steps. First, $N$ particles, i.e., realizations of the dynamic process, are propagated through the latent space over time. This means that, given the trajectory of the latent process for the $k$-th particle and the observables until time $t_{i-1}$, we sample the state of the latent process at time $t_{i}$, $X_{t_{i}}^{(k)}\sim q_{t_{i}}\left(\cdot\mid X_{t_{0}:t_{i-1}}^{(k)},y_{t_{0}:t_{i-1}},\theta\right)$, based on a chosen proposal distribution $q_{t_{j}}$. Second, at each time point, particles are weighted, based on their agreement with the observed values for that time point. And third, they are resampled using auxiliary variables $A_{t_{j}}^{1:N}$ to discard low-weight particles. The resampled particles are then again propagated forward to the next time point. Hence, for each time $t_{i}$, the particle filter constructs a system of $N$ weighted particles $\{X_{t_{i}}^{(k)},\,w_{t_{i}}^{(k)}\}_{k=1}^{N}$ that approximates the filtering distribution $p(X_{t_{i}}\,\mid\,y_{t_{0}:t_{i}},\theta)$. Mathematically, this is closely linked to the framework of Feynman-Kac models [moral2004feynman, andrieu_2010, chopin_2020].
@@
8:   Sample $A^{1:N}_{t_{i}}\sim\text{resample}(W_{t_{i}}^{1:N})$    $\triangleright$ See Supplementary Information A.4.2
@@
Although the name filtering refers to the task of estimating the latent state $X_{t_{i}}$ online given data $y_{t_{0}:t_{i}}$, a PF yields, as a by-product, a scheme to compute an unbiased estimator of the likelihood contributions ${\{p(y_{t_{0}:t_{i}}\,\mid\,\theta)\}}_{i\geq 0}$, crucial for the task of parameter inference [chopin_2020]. This estimator is obtained by computing the normalizing constants
@@
In this work, we use the likelihood estimate obtained by the bootstrap filter within the acceptance probability of a Metropolis-Hasting (MH) algorithm as the outer MCMC scheme, making it a member of the class of Pseudo-Marginal Metropolis-Hasting methods (PMMH) [andrieu_roberts, andrieu_2010, chopin_2020].
Using an unbiased estimator $\hat{p}(y|z,\theta)$ of the $p(y|\theta)$ that is obtained by the use of auxiliary random variables $z\sim r(\cdot)$ within the acceptance of a MH algorithm then corresponds to simulating a Markov chain $\{\theta^{i},z^{i}\}_{i\geq 0}$ targeting the joint density
@@
For the bootstrap filter, we can, for instance, think of the $A_{t_{j}}^{1:N}$ used in the resampling as being the auxiliary random variables. The joint density admits the correct posterior density of the model $p(\theta\mid y)$ as a marginal and therefore the algorithm samples the parameter vectors $\theta$ from the true posterior [andrieu_roberts, andrieu_2010, doucet_2015]. With respect to an extended distribution, PMMH algorithms are standard MCMC samplers.
One can use any unbiased estimator within any MCMC algorithm to produce a PMCMC algorithm. Different filtering and sampling algorithms impose different practical and implementational challenges and we focus here on the most general applicable version [doucet_2015, chopin_2020]. In the subsequent sections PF method always refers to a PMMH algorithm consisting of a bootstrap filter inside an adaptive Metropolis-Hastings algorithm.
@@ 3.2 Neural Posterior Estimation And Conditional Normalizing Flows
Neural posterior estimation (NPE) is a simulation-based inference (SBI) method which frames posterior construction as a (conditional) density-estimation problem. Given a simulator that produces samples $(\mathcal{D},\theta)$ by first drawing $\theta\sim p(\theta)$ and then $\mathcal{D}\sim p(\mathcal{D}\mid\theta)$, NPE fits a conditional neural network $p_{\phi}(\theta\mid\mathcal{D})$ to approximate the true posterior $p(\theta\mid\mathcal{D})$. Typically the data $\mathcal{D}$ are preprocessed by a (learned) summary statistic [radev2020bayesflow], which serves as the input to the conditional density estimator. For clarity, this preprocessing step is omitted from the notation below.
In this work, we focus on Conditional Normalizing Flows (CNFs) [rezende2015variational]. In general, normalizing flows transform a simple base distribution (e.g., a standard Gaussian) into a complex target distribution via a sequence of invertible, differentiable mappings. In our setting, this means that a normalizing flow defines a diffeomorphism $f_{\mathcal{D}^{*}}$ such that (for fixed $\mathcal{D}^{*}$) the posterior $p(\theta\mid\mathcal{D}^{*})$ is the “pushforward” of the base distribution $q$ along $f_{\mathcal{D}^{*}}$, i.e.,
@@
CNFs extend this idea by allowing the transformation to additionally depend on datasets $\mathcal{D}$, generated by first drawing $\theta\sim p(\theta)$ and then $\mathcal{D}\sim p(\mathcal{D}\mid\theta)$. Rather than learning a single transformation $f$, CNFs learn a family of conditional flows $\{f_{\mathcal{D}}\}$, with each $f_{\mathcal{D}}$ mapping the base distribution $q$ to the posterior $p(\theta\mid\mathcal{D})$ [winkler2019conditional]. This conditioning enables CNFs to capture how the posterior changes with different observations, making them particularly well-suited for amortized inference tasks, where a model learns to perform inference efficiently across multiple datasets rather than solving each case from scratch [kobyzev2020normalizing, rezende2015variational].
Instead of hand-designing each $f_{\mathcal{D}}$, we parameterize the family using a neural network $f_{\phi}(\cdot,\,\mathcal{D})$ whose parameters $\phi$ are shared across all $\mathcal{D}$, but whose behavior is modulated by $\mathcal{D}$ itself [ardizzone2019guided]. We train the neural network by minimizing the Kullback-Leibler divergence between the true and learned posteriors for all $\mathcal{D}$ in the training distribution [weilbach2020structured, siahkoohi2020conditional, radev2023bayesflow]. This means, our optimization objective is finding $\hat{\phi}$
@@
1:Data source $S$ (pre‑simulated dataset or simulator callable $\sim p(\theta)p(\mathcal{D}\mid\theta)$), base density $q(z)$, conditional flow $f_{\phi}(\cdot;\mathcal{D})$, optimizer Opt, learning rate $\eta$, batch size NN, iterations $T$, inference data $\mathcal{D}^{*}$, inference sample count $M$
2:Trained parameters $\phi$; sampler for approximate posterior $p_{\phi}(\theta\mid\mathcal{D}_{*})$
3:for $t=1$ to $T$ do
4:  $\{(\theta_{i},\mathcal{D}_{i})\}_{i=1}^{N}\leftarrow\text{sample from $S$ (dataset or simulator)}$
5:  for $i=1$ to NN do
6:   $z_{i}\leftarrow f_{\phi}(\theta_{i};\mathcal{D}_{i})$
7:   $\log p_{\phi}(\theta_{i}\mid\mathcal{D}_{i})\leftarrow\log q(z_{i})+\log\Big|\det\Big(\frac{\partial f_{\phi}(\theta_{i};\mathcal{D}_{i})}{\partial\theta}\Big)\Big|$
8:  end for
9:  $\mathcal{L}\leftarrow-\frac{1}{N}\sum_{i=1}^{N}\log p_{\phi}(\theta_{i}\mid\mathcal{D}_{i})$
10:  $\phi\leftarrow\text{Opt}(\phi,\nabla_{\phi}\mathcal{L},\eta)$
11:end for
12:
@@
14:  for $m=1$ to $M$ do
15:   Sample $z_{m}\sim q(z)$
16:   $\theta_{m}\leftarrow f_{\phi}^{-1}(z_{m};\mathcal{D}_{*})$
17:  end for
18:  return $\{\theta_{m}\}_{m=1}^{M}$
19:end procedure
This amortization not only enables fast inference once training is complete, but also facilitates simulation‑based calibration (SBC) diagnostics by allowing posterior approximations to be generated efficiently across many simulated datasets. At the same time, performance of CNFs is tied to the support of the training distribution: the model can interpolate to new values if they are close to the training support, but performance degrades when extrapolating far beyond it [winkler2019conditional, talts2018validating].

### v3

@@ 2.4 Observation Model
The course of an epidemic is usually only partially observed and data is often only available in aggregated form on population level. For all models introduced above, we consider an observation model reflecting one to two often measured quantities to monitor infectious diseases. For the SIS model, the observable is
$y(t)=\left(\begin{array}[]{c}\frac{\mathrm{I}_{\text{}}(t)}{\mathrm{N}_{\text{}}(t)}\end{array}\right).$ (3)
For the SIR model, the observables are
@@
where the first entry is the ratio of currently infected people describing the count data from testing, and the second entry describes seroprevalence, the ratio of individuals with antibodies present due to a prior infection. Because the observation process may have several limitations, an additional scaling parameter $s$ is used to capture that only a fraction of cases is reported [Gudina et al., 2021].
@@
where we truncate $\overline{y}_{1}(t_{i})$ and $\overline{y}_{2}(t_{i})$ at a lower bound of 0 to avoid unrealistic negative observations. As in the dataset given by Gudina et al. [2021], the standard deviations $\sigma_{i,1},\sigma_{i,2}$ were computed directly using a normal approximation to the binomially distributed testing count data. Alternative probability distributions may be more appropriate in other contexts.
From each of the stochastic compartmental models we generated synthetic data emulating the time course of a disease outbreak. We generated dense datasets with observations every seven or ten days. We used two types of parameter regimes:
  1. 1.
    hand‑selected values chosen to represent plausible epidemiological regimes, and
  2. 2.
    parameters sampled directly from the corresponding priors as an unbiased baseline
Prior draws that generated trajectories with collapsed variance were discarded and resampled to maintain numerical stability of the particle filter. Additionally, to create realistic, heterogeneous observation schedules, we randomly subsampled the dense time series into sparse datasets reducing the number of time points from 57 to fewer than 15 per observable. We also introduced partial missingness in the sparse datasets, where only one of the two observables was available at certain time points. Details on the computation of measurement noise and data generation are provided in the Models section of the Supplementary Information S1 (Section S1.3).
@@ 3 Bayesian Inference for Stochastic Compartmental Models
To infer the unknown model parameters $\mathbf{\theta}$ given data $\mathcal{D}^{*}=\{\overline{y}_{t_{i}}\mid i\in\{1,...,n\}\}$}, we apply Bayesian inference techniques. For this, we denote the space of unobserved states of the diffusion process by $\mathcal{X}$.
@@
The posterior represents the updated belief about the parameters after taking the observed data into account via the likelihood, effectively integrating both previous knowledge and the new information provided by the data. While the full posterior distribution is typically of primary interest, we additionally consider the mode (called maximum a-posteriori) of the posterior distribution as the best point estimate [Bernardo et al., 1994].
@@
In most cases, evaluating the density of hidden states analytically is impossible [Doucet et al., 2001], which prohibits evaluating the likelihood. Even in the case of a tractable likelihood, normalization of the posterior is done via the marginal likelihood, which can remain intractable. High‑dimensional parameter spaces further make discretization or numerical integration prohibitively expensive [Wilkinson, 2011; Hasenauer, 2013].
Approximation techniques that estimate the posterior distribution without directly computing the complicated likelihood function enable robust parameter estimation, even in settings where traditional methods may fail due to mathematical or computational complexity. In this section, we will introduce two approaches (particle filters (PF) [Golightly and Wilkinson, 2011] and one particular class of conditional density estimators, conditional normalizing flows (CNF) [Papamakarios and Murray, 2016; Cranmer et al., 2020]) that circumvent the computation of the likelihood function. PF approximate the likelihood with an unbiased estimator obtained by sequential importance sampling, and conditional normalizing flows learn a series of invertible mappings to transform a Gaussian density into the desired density.
@@ 3.1 Particle Markov Chain Monte Carlo Methods
Particle Markov chain Monte Carlo (PMCMC) methods are based on the idea of using an unbiased estimator of the likelihood function, obtained by a particle filter, inside a Markov chain Monte Carlo (MCMC) method [Golightly and Wilkinson, 2011].
Particle filters are sequential Monte Carlo (SMC) techniques based on importance sampling with resampling used for handling non-linear and non-Gaussian time-series models. Let $X_{t_{0}:t_{M}}$ denote a discrete time-series of the unobserved states and $y_{t_{0}:t_{M}}$ a time-series of observations at the same time points. The main idea is to iterate over the following three steps. First, $N$ particles, i.e., realizations of the dynamic process, are propagated through the latent space over time. This means that, given the trajectory of the latent process for the $k$-th particle and the observables until time $t_{i-1}$, we sample the state of the latent process at time $t_{i}$, $X_{t_{i}}^{(k)}\sim q_{t_{i}}\left(\cdot\mid X_{t_{0}:t_{i-1}}^{(k)},y_{t_{0}:t_{i-1}},\theta\right)$, based on a chosen proposal distribution $q_{t_{j}}$. Second, at each time point, particles are weighted, based on their agreement with the observed values for that time point. And third, they are resampled using auxiliary variables $A_{t_{j}}^{1:N}$ to discard low-weight particles. In practice, the resampling is not done in every iteration, but only if the variability of the weights becomes too large, as measured by the effective sample size (ESS) (for more details see Supplementary Information S1.1.2.2). The resampled particles are then again propagated forward to the next time point. Hence, for each time $t_{i}$, the particle filter constructs a system of $N$ weighted particles $\{X_{t_{i}}^{(k)},\,w_{t_{i}}^{(k)}\}_{k=1}^{N}$ that approximates the filtering distribution $p(X_{t_{i}}\,\mid\,y_{t_{0}:t_{i}},\theta)$. Mathematically, this is closely linked to the framework of Feynman-Kac models [Moral, 2004; Andrieu et al., 2010; Chopin and Papaspiliopoulos, 2020].
@@
8:   Sample $A^{1:N}_{t_{i}}\sim\text{resample}(W_{t_{i-1}}^{1:N})$    $\triangleright$ See Supplementary Information S1.1.2.2
@@
Although the name filtering refers to the task of estimating the latent state $X_{t_{i}}$ online given data $y_{t_{0}:t_{i}}$, a PF yields, as a by-product, a scheme to compute an unbiased estimator of the likelihood contributions ${\{p(y_{t_{0}:t_{i}}\,\mid\,\theta)\}}_{i\geq 0}$, crucial for the task of parameter inference [Chopin and Papaspiliopoulos, 2020]. This estimator is obtained by computing the normalizing constants
@@
In this work, we use the likelihood estimate obtained by the bootstrap filter within the acceptance probability of a Metropolis-Hasting (MH) algorithm as the outer MCMC scheme, making it a member of the class of Pseudo-Marginal Metropolis-Hasting methods (PMMH) [Andrieu and Roberts, 2009; Andrieu et al., 2010; Chopin and Papaspiliopoulos, 2020]. The full PMMH-algorithm is provided in the Supplementary Information S1.1.1.
Using an unbiased estimator $\hat{p}(y|z,\theta)$ of the likelihood $p(y|\theta)$ that is obtained by the use of auxiliary random variables $z\sim r(\cdot)$ within the acceptance of a MH algorithm then corresponds to simulating a Markov chain $\{\theta^{i},z^{i}\}_{i\geq 0}$ targeting the joint density
@@
For the bootstrap filter, we can, for instance, think of the $A_{t_{j}}^{1:N}$ used in the resampling as being the auxiliary random variables. The joint density admits the correct posterior density of the model $p(\theta\mid y)$ as a marginal and therefore the algorithm samples the parameter vectors $\theta$ from the true posterior [Andrieu and Roberts, 2009; Andrieu et al., 2010; Doucet et al., 2015]. With respect to an extended distribution, PMMH algorithms are standard MCMC samplers.
One can use any unbiased estimator within any MCMC algorithm to produce a PMCMC algorithm. Different filtering and sampling algorithms impose different practical and implementational challenges and we focus here on the most general applicable version [Doucet et al., 2015; Chopin and Papaspiliopoulos, 2020]. In the subsequent sections PF method always refers to a PMMH algorithm consisting of a bootstrap filter inside an adaptive Metropolis-Hastings algorithm.
@@ 3.2 Neural Posterior Estimation And Conditional Normalizing Flows
Neural posterior estimation (NPE) is a simulation-based inference (SBI) method which frames posterior construction as a (conditional) density-estimation problem [Papamakarios and Murray, 2016]. Given a simulator that produces samples $(\mathcal{D},\theta)$ by first drawing $\theta\sim p(\theta)$ and then $\mathcal{D}\sim p(\mathcal{D}\mid\theta)$, NPE fits a conditional neural network $p_{\phi}(\theta\mid\mathcal{D})$ to approximate the true posterior $p(\theta\mid\mathcal{D})$. Typically the data $\mathcal{D}$ are preprocessed by a (learned) summary statistic [Radev et al., 2020], which serves as the input to the conditional density estimator. For clarity, this preprocessing step is omitted from the notation below.
In this work, we focus on Conditional Normalizing Flows (CNFs) [Rezende and Mohamed, 2015]. Normalizing flows have since been substantially developed and systematized [Papamakarios et al., 2021; Boelts et al., 2025; Radev et al., 2023]. In general, normalizing flows transform a simple base distribution (e.g., a standard Gaussian) into a complex target distribution via a sequence of invertible, differentiable mappings. In our setting, this means that a normalizing flow defines a diffeomorphism $f_{\mathcal{D}^{*}}$ such that (for fixed $\mathcal{D}^{*}$) the posterior $p(\theta\mid\mathcal{D}^{*})$ is the “pushforward” of the base distribution $q$ along $f_{\mathcal{D}^{*}}$, i.e.,
@@
CNFs extend this idea by allowing the transformation to additionally depend on datasets $\mathcal{D}$, generated by first drawing $\theta\sim p(\theta)$ and then $\mathcal{D}\sim p(\mathcal{D}\mid\theta)$. Rather than learning a single transformation $f$, CNFs learn a family of conditional flows $\{f_{\mathcal{D}}\}$, with each $f_{\mathcal{D}}$ mapping the base distribution $q$ to the posterior $p(\theta\mid\mathcal{D})$ [Winkler et al., 2019]. This conditioning enables CNFs to capture how the posterior changes with different observations, making them particularly well-suited for amortized inference tasks, where a model learns to perform inference efficiently across multiple datasets rather than solving each case from scratch [Kobyzev et al., 2020; Rezende and Mohamed, 2015].
Instead of hand-designing each $f_{\mathcal{D}}$, we parameterize the family using a neural network $f_{\phi}(\cdot,\,\mathcal{D})$ whose parameters $\phi$ are shared across all $\mathcal{D}$, but whose behavior is modulated by $\mathcal{D}$ itself [Ardizzone et al., 2019]. We train the neural network by minimizing the Kullback-Leibler divergence between the true and learned posteriors for all $\mathcal{D}$ in the training distribution [Weilbach et al., 2020; Siahkoohi et al., 2020; Radev et al., 2023]. This means, our optimization objective is finding $\hat{\phi}$
@@
1:Data source $S$ (pre‑simulated dataset or simulator callable $\sim p(\theta)p(\mathcal{D}\mid\theta)$), base density $q(z)$, conditional flow $f_{\phi}(\cdot;\mathcal{D})$, optimizer Opt, learning rate $\eta$, batch size NN, inference data $\mathcal{D}^{*}$, inference sample count $M$
2:Trained parameters $\phi$; sampler for approximate posterior $p_{\phi}(\theta\mid\mathcal{D}_{*})$ repeat
@@
4:for $i=1$ to NN do
5:  $z_{i}\leftarrow f_{\phi}(\theta_{i};\mathcal{D}_{i})$
6:  $\log p_{\phi}(\theta_{i}\mid\mathcal{D}_{i})\leftarrow\log q(z_{i})+\log\Big|\det\Big(\frac{\partial f_{\phi}(\theta_{i};\mathcal{D}_{i})}{\partial\theta}\Big)\Big|$
7:end for
[…]
10:
[…]
12:  for $m=1$ to $M$ do
13:   Sample $z_{m}\sim q(z)$
14:   $\theta_{m}\leftarrow f_{\phi}^{-1}(z_{m};\mathcal{D}_{*})$
15:  end for
16:  return $\{\theta_{m}\}_{m=1}^{M}$
17:end procedure
This amortization not only enables fast inference once training is complete, but also facilitates simulation‑based calibration (SBC) diagnostics by allowing posterior approximations to be generated efficiently across many simulated datasets. At the same time, performance of CNFs is tied to the support of the training distribution: the model can interpolate to new values if they are close to the training support, but performance degrades when extrapolating far beyond it [Winkler et al., 2019; Talts et al., 2018]. In practice, this amortized structure is precisely what makes SBC feasible at scale, since once trained, the CNF can be evaluated repeatedly on equally structured datasets without retraining.